# Unified Medical Image Pre-training in Language-Guided Common Semantic Space

### v1

@@ 4.3.1 Results on Medical Classification
3D Medical Volume Classification  Table 3 reports the results of Linear Classification on the 2D medical image classification dataset, CC-CCII. We compare UniMedI with UniMiss Xie et al. (2022). To our knowledge, the UniMiSS Xie et al. (2022) is the state-of-the-art unified method to process 2D and 3D medical images. We show the performances of both UniMiSS and UniMedI, where the results are that our method achieves a +22.6%, +2.0% and +0.8% ACC gain on CC-CCII dataset comparing with the UniMiSS under the 1%, 10%, 100% training ratio respectively. The significant improvement indicates the data efficiency and effectiveness of our method.
When fine-tuning the total vision encoder and the linear classification head with full training data, as listed in Table 3, our method gets the best performance on the multiple 3D medical volume classification datasets (CC-CCII and LUNA2016-v2) compared with other methods. It is observed that our method achieves with 93.8% ACC on CC-CCII dataset, and 95.9% ACC on LUNA2016-v2 dataset respectively. The remarkable performance of our method shows the generalization of our method for 2D and 3D medical classification tasks. It demonstrates our framework possesses the ability of extracting universal features for multi-modal data.
@@
Table 6: 2D Semantic segmentation results (Dice [%]) on RSNA with 1%, 10% and 100% training labels. Best results of each setting are in boldface.
@@
UniMedI (ViT-B/16) 67.8 73.1 75.3
@@
Table 6: 2D Semantic segmentation results (Dice [%]) on RSNA with 1%, 10% and 100% training labels. Best results of each setting are in boldface.
4.3.2 Results on Medical Semantic Segmentation
Table 7 and Table 7 report the results of Semantic Segmentaion on 2D and 3D medical data. In 2D semantic segmentation task, our method UniMedI significantly outperforms the current state-of-the-art algorithm, MGCA. When using 1% training data, UniMedI achieves 67.8% Dice, surpasssing the MGCA 1.6%. Meanwhile, concurrently, UniMedI also demonstrates exceptional performance in 3D semantic segmentation tasks. In the BCV dataset, UniMedI achieves 0.6% and 0.4% performance gain under 20% and 40% label settings compared with Unimiss. These results underscore the exceptional performance of our method in dense prediction tasks.
4.4 Analysis of Our Framework
Figure 5: Visualization of mask and slices selection result under the guidance of language.
@@ 4.4 Analysis of Our Framework
Visualization To better demonstrate the effectiveness of our selection process guided by language, we visualize the original X-rays, masked X-rays, their corresponding reports, and original CT scans, as well as the selected lesion slices in Figure 5. On the left side of Figure 5, the first row effectively demonstrates how UniMedI accurately captures the areas referenced in the report, including the “Normal post-operative alignment of the sternal wires” and “Bilateral pleural effusions of mild-to-moderate extent persist”. In addition, the second and third cases adeptly showcase the detection of pleural effusion and scoliosis, further emphasizing the method’s precision. The right side of Figure 5 displays the comprehensive slice selection process employed by UniMedI. Amidst the extensive collection of CT scan slices, our method exhibits remarkable accuracy in pinpointing the slices containing lesions. As an example, the presence of pulmonary nodules is clearly noticeable in slices 28-31.
Ablation Study of Component Design We conduct ablation experiments primarily focusing on two aspects: training mode and framework module.
Training mode  We pre-train our framework separately using only 2D data, only 3D data, and a combination of 2D and 3D data. Subsequently, we evaluated the performance on downstream 2D dataset CheXpert, RSNA and 3D dataset CC-CCII on linear classification task respectively, with the results presented in Table 3. It can be observed that the pretraining approach combining 2D and 3D data yields benefits for both single-modal 2D and 3D data classification tasks. Particularly, the enhancement achieved with the use of multimodal data on the 3D dataset is remarkably significant. We obtained improvements of +16.8% ACC, +8.3% ACC, +9.8% ACC when using 1%, 10%, and 100% of the training data, respectively.
@@
Framework module  In this section, we further analyze the effects of self feature distillation and attentive slices selection on our framework. We conduct a linear classification task on downstream 2D datasets CheXpert and RSNA, as well as the 3D dataset CC-CCII. The results are summarized in Table 5. The experimental results show that incorporating both self feature distillation and attentive slices selection into our framework significantly improves the performance across all data splits and datasets.
@@ 5 Conclusion
Appendix A More Implementation Details of Pre-Training
A.1 Implementation Details
Following Gloria Huang et al. (2021), we utilize ViT-B/16 Dosovitskiy et al. (2020) as the vision encoder to extract representations in the common feature space for 2D and 3D visual data. We use BioClinicalBERT Alsentzer et al. (2019) as the text encoder to obtain the report embeddings. The vision encoder and text encoder are universal among 2D X-rays and 3D CT scans data. It is worth noting that the patch embed module of vision encoder has different operations for 2D X-rays and 3D CT scans. In general, the image size of 2D images is $224\times 224$ and the volume size of 3D volumes is $128\times 128\times 32$. We pre-train our UniMedI framework 50 epochs on 8 pieces of Tesla V100 GPUs with batch size of 128. The optimizer is AdamW Loshchilov & Hutter (2017) with learning rate of $2e^{-5}$ and weight decay of 0.05, where the learning rate follows a linear warmup with cosine annealing scheduler Loshchilov & Hutter (2016). We initialize learning rate as $1e^{-8}$ and warmup epoch as 20.
Appendix B More Implementation Details of Downstream Tasks
B.1 Medical Classification
2D Medical Image Classification. Except for the fine-tuning of the entire CheXpert dataset, where we use a batch size of 96, we use a batch size of 48 for the rest of the linear classification settings. Similar to the image preprocessing of MIMIC-CXR, we resize the larger dimension to 256 and pad zeros on the smaller side, resulting in an image size of 256 $\times$ 256. Then, we randomly crop (for training) or centered crop (for validation and testing) an image to 224 $\times$ 224 and normalize it into the range [0, 1] as the input for the model. The optimizer used is AdamW Loshchilov & Hutter (2017) with a learning rate of 5e-4 (except for COVIDx where we use 5e-3) and weight decay of 1e-6. We fine-tune the image classifier for 50 epochs and implement early stopping when the validation loss does not decrease for 10 consecutive runs. Afterward, we save the checkpoint model with the lowest validation loss for testing.
3D Medical Image Classification. (1) CC-CCII Zhang et al. (2020) contains a total number 617,775 slices of 6,752 CT scans from 4,154 patients. The task is to classify each volume into three categories: novel coronavirus pneumonia, common pneumonia, and normal. We use a batch size of 8. We resize the 3D volumes to $32\times 128\times 128$. We use the randomflip to augment the train set. The optimier used is AdamW and we train the classifier for 50 epochs. (2) LUNA 16 Setio et al. (2017), which is established from LIDC-IDRI Armato III et al. (2011). It finally contains 888 CT scans with annotations which removes CT scans with slice thickness greater than 3mm of LIDC-IDRI database. The task is a binary classification, i.e., classifying each CT volume into pulmonary nodule or normal. The optimier used is AdamW and we train the whole network for 100 epochs. Our baseline methods include UniMiss Xie et al. (2022) and Joint Nguyen et al. (2023), which belongs to 2D and 3D co-learning methods. Unimiss not only learns 2D and 3D representations but also concurrently learn all 2D sections derived from 3D volumes, along with all 2D X-ray data. Joint directly learn all 2D sections derived from 3D volumes, along with all 2D X-ray data through contrastive learning.
B.2 Medical Segmentation.
2D Medical Image Segmentation. In the case of the RSNA dataset, we create masks for the pneumonia-affected areas based on the provided bounding boxes. These images and corresponding masks are then resized to dimensions of 224 $\times$ 224. To augment the training set, we implement ShiftScaleRotate, encompassing random affine transformations such as translation, scaling, and rotation. Following this, the images are normalized to fall within the [0, 1] range before being supplied to the semantic segmentation model. we use the SETR-PUP (progressive upsample) architecture in Zheng et al. (2021) by replacing the encoder with UniMedI. We freeze the pre-trained image encoder and only train decoder portion. The training process involves the use of the AdamW optimizer with a learning rate of 5e-4 and a weight decay of 1e-6. As suggested by Huang et al. (2021), we adopt a combined loss equation of $\alpha\times$ FocalLoss + DiceLoss, with $\alpha$ set to 10. The semantic segmentation model undergoes fine-tuning for 50 epochs, with batch size 16 and early stopping implemented if the validation loss ceases to decrease after 10 consecutive runs. The checkpoint model that exhibits the minimum validation loss is then preserved for testing.
3D Medical Image Segmentation. In the case of the BCV dataset, the images and correspoinding. The 3D volumes are resized to $48\times 224\times 224$. To augment the traning set, we implement random rotation, scaling, flipping, adding white Gaussian noise, Gaussian blurring, adjusting rightness and contrast, simulation of low resolution, and Gamma transformation. We use the UNETR Hatamizadeh et al. (2022) architecture by replace the encoder with pre-trained UniMedI. We freeze the pre-trained image encoder and only train decoder portion. The training process involves the use of the AdamW optimizer with a learning rate of 1e-4. We adopt a combined loss equation of Dice + CE. The semantic segmentation model fintunes for 25,000 iterations with batch size 2.
Appendix C More Analysis
C.1 Pneumonia Detection in RSNA
We evaluate the localized performance of pre-trained image encoder on RNSA Pneumonia. RSNA contains 29700 frontal view radiograph. The task is to predict bounding boxes indicating evidence of pneumonia. Due to use ViT-B as our bakcbone, it is sufficient to build a simple feature pyramid from a single-scale feature map. Therefore, we evaluate the detection performance by ViTDet Li et al. (2022) with using the pre-trained ViT-B as a frozen backbone and only finetuning the non-backbone layers. Similarly, we finetune the model by 1%, 10% and 100% training data to evaluate the data efficiency.
RSNA
Method 1% 10% 100%
ConVIRT 8.2 5.6 17.9
GLoRIA 9.8 14.8 18.8
GLoRIA-MIMIC 11.6 16.1 24.8
MGCA (ResNet-50) 12.9 16.8 24.9
MGCA (ViT-B) 14.7 18.4 25.8
UniMedI 15.5 19.2 26.6
Table 8: Object detection results (mAP [%]) on RSNA. Each dataset is fine-tuned with 1%, 10%, 100% training data. Best results are in boldface.
C.2 3D Medical Segmentation in AMOS
AMOS is a large-scale, diverse, clinical dataset for abdominal organ segmentation, which is divided into 200/100 CTs for training/validation. We use the validation set as our test set and the training details is the same as B.2. We report the Dice score (%) training with 20%, 40%, and 100% portion.
AMOS
Method 20% 40% 100%
UniMiss 79.5 82.3 85.8
UniMedI 78.8 82.9 86.4
Table 9: 3D Semantic Segmentation results on AMOS (Dice [%]). AMOS is fine-tuned with 20%, 40%, 100% training data. Best results are in boldface.
C.3 Different Metrics in COVIDx
COVIDx (Acc/AUC)
Method 1% 10% 100%
MGCA 74.8/89.0 84.8/97.0 92.3/97.9
UniMedI 80.3/93.5 92.4/98.1 94.6/98.1
Table 10: 3D Semantic Segmentation results on AMOS (Dice [%]). AMOS is fine-tuned with 20%, 40%, 100% training data. Best results are in boldface.
We applied two distinct evaluation metrics, namely AUC (Area Under the Curve) and ACC (Accuracy), to assess the performance of our model on the COVIDx dataset. AUC is a widely used metric in machine learning and it represents the probability that a random positive example will be ranked higher than a random negative example. A higher AUC indicates better model performance. On the other hand, Accuracy (ACC) is a measure of how many predictions a model gets right out of all the predictions it makes. It is calculated as the number of correct predictions divided by the total number of predictions. The results of our evaluation using these metrics on the COVIDx dataset are presented in Table 10. These findings provide insights into the robustness of our model.
C.4 Visualization of Feature Quality
We add three t-SNE visualization in Figure. 6. Compared to Figure. 2, Figure. 6 add more class (cardiomegaly) to demonstrate the ability to unify different modal representations. We have marked the changes in distances between different modalities in the figure. As shown in Figure. 6, UniMedI effectively reduces the distance between different modalities and stablishes a unified representation space.
Figure 6: t-SNE visualizations of image representations by models trained with different methods. (a) Two models for different image modalities are trained individually in separate VLP process. (b) One models for different image modalities are trained in one VLP processes, but without designes in UniMedI. (c) UniMedI. We use circles to highlight differences between different images.
=" alt="[LOGO]">

### v2

@@ 4.4.2 Results on 3D Medical Volume Classification
Table 3 reports the results of linear classification on the 3D medical image classification dataset, CC-CCII. We compare UniMedI with UniMiss [36]. To our knowledge, the UniMiSS [36] is the state-of-the-art unified method to process 2D and 3D medical images. We show the performances of both UniMiSS and our UniMedI, where the results are that our method achieves a +22.6%, +2.0% and +0.8% ACC gain on CC-CCII dataset comparing with the UniMiSS under the 1%, 10%, 100% training ratio respectively. The significant improvement indicates the data efficiency and effectiveness of our method.
@@
Table 4: 2D Semantic segmentation results (Dice [%]) on RSNA with 1%, 10% and 100% training labels.
@@
UniMedI 67.8 73.1 75.3
Table 5: 2D Object detection results (mAP [%]) on RSNA with 1%, 10%, 100% training data.
RSNA
Method 1% 10% 100%
ConVIRT 8.2 5.6 17.9
GLoRIA 9.8 14.8 18.8
GLoRIA-MIMIC 11.6 16.1 24.8
MGCA 12.9 16.8 24.9
MGCA (ViT-B/16) 14.7 18.4 25.8
UniMedI 15.5 19.2 26.6
When fine-tuning the total vision encoder with full training data, as listed in Table 3, our method gets the best performance on the multiple 3D medical volume classification datasets (CC-CCII and LUNA) compared with other methods. It is observed that our method achieves with 93.8% ACC on CC-CCII dataset, and 95.9% ACC on LUNA2016-v2 dataset respectively. The remarkable performance of our method shows the generalization of our method for 2D and 3D medical classification tasks. It demonstrates our framework possesses the ability of extracting universal features for 3D CT scans.
4.4.3 Results on 2D Medical Semantic Segmentation
We evaluate the impact of local representation by analyzing its performance in the task of semantic segmentation. Table 5 presents the results of all methods on the RSNA dataset. UniMedI significantly outperforms the current state-of-the-art algorithm, MGCA. When using 1% training data, UniMedI achieves 67.8% Dice, surpasssing the MGCA 1.6%. In addition to segmentation tasks, we also conducted detection tasks on the RSNA Pneumonia dataset to demonstrate the localized capabilities of UniMedI. The objective is to identify and predict bounding boxes that signify the presence of pneumonia. Utilizing ViT-B as our backbone, we can create an effective feature pyramid from a single-scale feature map, which simplifies the architecture. For the detection task, we leveraged ViTDet [20] with a pre-trained ViT-B serving as a frozen backbone, while only the non-backbone layers were fine-tuned. To assess data efficiency, we fine-tuned the model with varying amounts of training data: 1%, 10%, and 100%. Table 5 presents the results of all methods on the RSNA dataset.
4.4.4 Results on 3D Medical Semantic Segmentation
Table 6 showcases the performance of UniMedI on the BCV dataset for semantic segmentation in 3D medical image. Compared to the baseline method UniMiSS, our approach yields a 0.6% and 0.4% increase in accuracy with limited label availability of 40% and 100%, respectively. These results underscore the superior capability of UniMedI in extracting meaningful features and effectively capitalizing on limited annotated data. This performance gain reflects UniMedI’s adeptness at harnessing local representations for the semantic segmentation task.
4.5 Visualization and Ablation Study
Figure 5: Visualization of masking and slices selection result under the guidance of language.
@@ 4.5 Visualization and Ablation Study
4.5.1 Visualization
@@
To enhance the clarity and showcase the efficacy of our linguistically-informed selection process, we conduct a comprehensive visual presentation in Fig. 5. This includes the display of the original X-ray images alongside the masked X-rays, where the regions of interest are prominently highlighted. Accompanying these images are the corresponding medical reports that provide context and diagnostic insights. Furthermore, we exhibit the original CT scans , as well as selected lesion slices. On the left side of Fig. 5, the first row effectively demonstrates how UniMedI accurately captures the areas referenced in the report, including the “Normal post-operative alignment of the sternal wires” and “Bilateral pleural effusions of mild-to-moderate extent persist”. In addition, the second and third cases adeptly showcase the detection of pleural effusion and scoliosis, further emphasizing the method’s precision. The right side of Fig. 5 displays the comprehensive slice selection process employed by UniMedI. Amidst the extensive collection of CT scan slices, our method exhibits remarkable accuracy in pinpointing the slices containing lesions. As an example, the presence of pulmonary nodules is clearly noticeable in slices 28-31.
(a) CheXpert Dataset.
(b) RSNA Dataset.
(c) RSNA Dataset.
(d) CC-CCII Dataset.
Figure 6: Comparative study of common semantic space versus separate. The results presented here are set in the context of linear classification (2D dataset CheXpert, RSNA and 3D dataset CC-CCII) settings. We report Area under ROC curve (AUROC [%]) on (a) CheXpert Classification and (b) RSNA Classification, Dice [%] on (c) RSNA Segmentation and Acc [%] on (d) CC-CCII Classification.
4.5.2 Importance of the Common Semantic Space
As illustrated in Fig 2, we provide visualizations for three types of feature spaces to elaborate on the motivation behind UniMedI. We now proceed to its quantitative analysis. As shown in Fig 6(d), conducting joint training on both modalities (2D $+$ 3D) within the same model shows certain improvements over single-modality training, but the enhancement is not significant. The design of UniMedI, through techniques like language-guided pseudo pairs and self distillation, enables better joint modeling of 2D and 3D information. This leads to substantial performance improvements across all settings, which cannot be achieved by simple adding 2D and 3D data together.
Table 7: Ablation study of component design. We report Area under ROC curve (AUROC [%]) on CheXpert and RSNA datasets, and (Acc [%]) on CC-CCII dataset. $VL$ represents the default experiment setting include image-text contrastive loss $L_{vl}$ with random slices selection. $SD$ will include $L_{icl}$ and $L_{pcl}$ loss to execute Self-Distillation. $Attn$ will use attentive slices selection instead of random. Best results of each setting are in boldface.
Training tasks CheXpert (AUC) RSNA (AUC) CC-CCII (Acc)
$VL$ $SD$ $Attn$ $1\%$ $10\%$ $100\%$ $1\%$ $10\%$ $100\%$ $1\%$ $10\%$ $100\%$
87.4 88.1 88.5 88.9 89.3 90.6 72.4 80.0 86.2
89.0 89.3 90.1 89.5 90.1 91.2 74.6 80.9 86.7
89.4 89.7 90.5 90.0 90.4 91.5 75.6 84.8 89.4
4.5.3 Ablation on Self-Distillation and Attentive Slices Selection
@@
Results in Table 7 show that incorporating both self distillation and attentive slices selection into our framework significantly improves the performance across all data splits and datasets, indicating the importance of key designs in UniMedI.